The patient weighs 22Kg and has an order for 190 mcg of medication to be infused in 53 mL of IV solution. The drop rate is 19gtts/min with tubing delivering 10gtts/mL. How many mcg/kg/min will the patient receive?

Please show work with formula.
Answer #1

53 ml=190mcg

1 ml= 190/53 mcg=10 gtt

But drop rate is 19gtt/ min

as 10 gtt=190/53mcg

therefore 19gtt=190/530 *19= 6.8mcg

so drop rate is 6.8 mcg/ min for 22 kg

in other words, 6.8/22=0.31mcg/kg/min patient is receiving .